A report on the literacy rate for a sample of countries revealed the following percentages: Country Literacy Rate (in percent) Turkey 55 Tunisia 32 Tonga 93 Benin 20 Bangladesh 25 Colombia 47 Iceland 99 Iran 37 a. Find the mean percentage b. Find the median percentage


1
Expert's answer
2021-11-15T16:47:04-0500

Arranged array: 20, 25, 32, 37, 47, 55, 93, 99.


a. Mean: "\mu=\frac{20+25+32+37+47+55+93+99}{8}=\frac{408}{8}=51."


b. Median: "M=\frac{37+47}{2}=42."